OpenStudy (anonymous):

which graph best represents the solution to the following system {-3x+y<1 {-x+y<=1

OpenStudy (anonymous):

a and b

OpenStudy (anonymous):

c and d

HanAkoSolo (jamierox4ev3r):

Okay so first of all, do you know how to go about solving the system of equations @cutegirl ? awww you just went offline.. but here, I'll give you a quick overview

HanAkoSolo (jamierox4ev3r):

so we have here: -3x+y<1 , and -x+y\(\le\)1 we have to put both equations into y=mx+b form. So they would look like this: y<3x+1 y\(\le\)x+1 these are your two equations. Then, graph them both on a graph. Here are some pointers: \(\checkmark\)equations with \(\gt\) or \(\lt\) are graphed with a dotted line \(\checkmark\)equations with \(\ge\) or \(\le\) are graphed with a straight-line \(\checkmark\)graph the two equations like normal equations… just be sure to note whether its less than or greater than. In this case, its less than (this is important so you can know where to shade). God I hope I'm making sense so far xD lol

HanAkoSolo (jamierox4ev3r):

 annnd I apologize for the really bad shading